Mpho has recently inherited a trust fund that offers two alternatives: she can take a R1 million pay-out now or receive an annual payment of R145 000 at the end of each year for 10 years. If she invests the money, she can receive 7% interest, compounded annually. Evaluate every option and advise her on which is the better option to take.




Your friend Fred is looking to buy a car and has a deposit of R50 000. The interest rate on offer is a fixed interest rate for 60 months of 7,5% per annum, compounded monthly. The car costs R300 000. What will the monthly repayment be?




Siya and Rachel want to ensure that they are able to fund their two children’s tertiary studies and have asked you to assist them in determining how much they need to save monthly to have R2 million in 18 years from now. They believe they can get an annual interest rate of 5%, compounded monthly. How much should they contribute to the fund monthly?


1
Expert's answer
2021-03-12T07:27:46-0500

1.If you put $1m into account, then you will get: "FV = 1*1.07^{10} =" $1.967m.

If use use annuity, then: FV = 145,000*10 = $1,450,000.

So, 1 is the better alternative.

2.The monthly payment is "A = \frac{250000*0.075/12*(1+0.075/12)^{60}}{(1+0.075/12)^{60}-1} = 5,009.49."

3."Pmt = \frac{2000000*0.05/12}{(1+0.05/12)^{18*12}-1} = 5,727.34."